OpenStudy (anonymous):

Systems of equations. 4x+3y=5, 2x-5y=-17?

OpenStudy (anonymous):

If you subtract 2 times the second equation from the first you will get an equation you can solve for y.

OpenStudy (anonymous):

Does that make sense or do you need further clarification?

OpenStudy (anonymous):

No, that makes sense. Thankyou.(:

OpenStudy (anonymous):

Wait, no it doesnt.. Im confused. Subtract what ?

OpenStudy (amistre64):

there are a few ways to solve systems of equation; one way was by substitution which can be rather labor intensive; the other method is called elimination, and that is what polpak suggests

OpenStudy (anonymous):

Okay... How do you eliminate?

OpenStudy (amistre64):

4x+3y=5 2x-5y=-17 <-- if we multiply this equation by -2 we get these results 4x+ 3y = 5 -4x+10y =34 now when we add these together, the x parts cancel out ------------ 0 +13y = 39 ; now we solve for y by dividing out the 13 to get: y = 39/13

OpenStudy (amistre64):

we can do the same thing to eliminate the y parts first, but it takes modifying both equations to do so; but still might be worth it to avoid dealing with fractions :)

OpenStudy (anonymous):

I ... how... blah. I need to do this in fractional form. How do you do that?

OpenStudy (anonymous):

Fractional form? I think that just means leave it as 39/13 instead of writing out the decimal answers.

OpenStudy (anonymous):

But what about x? Thats only the answer for y, rght?

OpenStudy (amistre64):

:) 4x +3y =5 <-- multiply by 5 2x -5y =-17 <-- multiply by 3 20x +15y =25 6x -15y =-51 ; now add them together ------------- 26x + 0 = -26 ; divide out the 26 to get: x = -26/26 = -1 ... if i did it right :)

OpenStudy (anonymous):

... I hate math. :l Why do you multiply the whole equation?

OpenStudy (anonymous):

Once you get an answer for y you can plug it into either equation to find x.

OpenStudy (anonymous):

You multiply the whole equation because you want to get the coefficients ( numbers in front of the variables) to match. In order to do something to one side of the equation (multiply by 3 for example) you must do the same thing to the other side so they will stay equal.

OpenStudy (amistre64):

to dbl check lets use these in the equations and solve 4x+3y=5 4(-1) + 3(39/13) = 5(13)/13 -4(13/13) + 117/13 = 65/13 117 - 52 = 65 .... its good there 2x-5y=-17 2(-1)-5(39/13)=-17(13)/13 -2(13)/13 -195/13 = -221/13 -26-195 = -221 .... and its good

OpenStudy (anonymous):

If I tell you that 5x + 3y = 17 Then 5 times 5x + 3y must equal 5 times 17.

OpenStudy (amistre64):

equations can change form while they keep the same value so what we want to do is change the form into something we can use

OpenStudy (anonymous):

... Okay. Hold on, let me work this all out and see if i get confused.

OpenStudy (amistre64):

:) i almost got confused on this one lol alot of fractional work to keep track of

OpenStudy (anonymous):

Certainly. What you're trying to do is find something you can multiply one (or both) of the equations by so that when you add the equations together you make one of the variables have a 0 for a coefficient.

OpenStudy (anonymous):

Alright! I did it and it all made sense and i got the same answers you guys did. (: Thank you.